Shift report 20030117 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Jan 17 2003 - 16:29:32 EST

  • Next message: Apache: "Shift report 20030118 16:00-24:00"
    ------------------------------------------------------------
    Date: Fri Jan 17 16:29:31 2003, Shift: 08:00-16:00
    Supervisor: or, with: 
    Log book#13 pages 77 - 78
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam until 15.30 pm.
    2. Start run 7435, FS 1/2 B at 4 degree, MRS 500 B at 90. We observe that trigger 6 not working and we stop the run at 15.57.
       Start the run 7436 to check trigger 6 but still not working. Stop the run at 16.13 because the beam is lost.
    3. H1 works now.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Jan 18 2003 - 14:47:03 EST